How do you do 1/4 is greater than or equal to 1/10q. I will need this by 10:00 please because I have to be in bed by 10:00.
Thanks:)

Answer by Alan3354(69443)

How do you do 1/4 is greater than or equal to 1/10q.
------------
Define "do"
----
If you want to solve for q, you should say that.
---
If you mean
1/4 >= (1/10)q
q/10 <= 1/4
Multiply by 10
q <= 5/2
=============
If you mean 1/4 >= 1/(10q) it's a different problem.
------------
1000 in what time zone?
It's 910 here in CT, 710 in CA, 510 in Hawaii